Question 9 (1 point) Instruction 3-1 Use the following tax rates, ceiling and maximum taxes: Employee and Employer OASDI: 6.20% Employee* and Employer HI: 1.45% Self-employed OASDI Self-employed HI: $128,400 No limit $128,400 No limit $7,960.80 No maximum $15,921.60 12.4% 2.9% No maximum *Employee HI: Plus an additional 0.9% on wages over $200,000. Also applicable to self-employed. Rounding Rules: Unless instructed otherwise compute hourly rate and overtime rates as follows: 1. Carry the hourly rate and the overtime rate to 3 decimal places and then round off to 2 decimals places (round the hourly rate to 2 decimal places before multiplying by one and one-half to determine the overtime rate) 2. If the third decimal place is 5 or more, round to the next higher cent. 3. If the third decimal place is less than 5, drop the third decimal place Also, use the minimum hourly wage of $7.25 in solving these problems and all that follow.
Employee HI: Plus an additional 0.9% on wages over $200,000. Also applicable to self-employed. Rounding Rules: Unless instructed otherwise compute hourly rate and overtime rates as follows: 1. Carry the hourly rate and the overtime rate to 3 decimal places and then round off to 2 decimals places (round the hourly rate to 2 decimal places before multiplying by one and one-half to determine the overtime rate) 2. If the third decimal place is 5 or more, round to the next higher cent 3. If the third decimal place is less than 5, drop the third decimal place Also, use th minimum hourty wage of $7 25 in solving these problens and all hat follow Refer to Instruction 3-1. Crow earned $585.15 during the week ended March 1, 20--. Prior to payday, Crow had cumulative gross earnings of $4,733.20. a) b) The amount of OASDI taxes to withhold from Crow's pay is The amount of HI taxes to withhold from Crow's pay is
Answer #1

Employee OASDI = 1,28,400

TAX rate as per self employed will be = 12.4%

Tax will be= 14,850.78

a) crow earned 585.15 during the week. Total yearly earning will be = 585.15*52= $30,427.80

Cumulative gross earnings is $4733.20

The amount of OASDI tax will be = $30,427.80+$4733.20* 12.4%= $4778.

b) the amount of HI tax will be = $30427.80+$4733.20*1.45%= $966.70